OpenStudy (anonymous):

2y-6/55 divided by y-3/77y divide and simplify

OpenStudy (anonymous):

\[\huge \frac{\frac{2y-6}{55}}{\frac{y-3}{77y}} \] factor the 2y-6 to 2(y-3) cancel the (y-3) factors. cancel an 11 from the 55 and 77. can you take it from here?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):